# Dynamical arrest in active nematic turbulence

### v2

@@ Abstract
A node with two incoming walls and one outgoing wall is only possible by combining different wall types (Supplementary Fig. 3). We further note that startpoints are very common in the non-arrested (contractile aligning) regime but become increasingly rare over time in the arrested (extensile aligning) regime.
We define these three types of points as pseudo-defects. Unlike proper topological defects, they are not phase singularities of the director field, and they do not carry a topological charge. Nevertheless, the director varies around them in a manner that resembles true nematic defects, except for its sharp counter variation across the walls. We thus assign the nodes a pseudo-charge, defined as the number of turns of the director along a path that surrounds the pseudo-defect, excluding the sharp variation across all intersecting walls (Figs. 3, 3 and 3). The excluded walls in this definition are analogous to the so-called Dirac strings, which are constructs that enable magnetic monopoles in Maxwell’s equations48 (see discussion in Supplementary Note). With this definition, endpoints have a positive pseudo-charge $+Q$ (Fig. 3) and both startpoints and branchpoints have pseudo-charge $-Q$ (Figs. 3 and 3). The pseudo-charge depends on $\nu$. For $S\nu>-1$, we have $Q\simeq 1/2$, while for $S\nu\leq-1$, we have $Q\simeq\frac{1}{2}\left[1-\frac{1}{\pi}\arccos{(|\nu|^{-1}})\right]$ (see details in Supplementary Note). The total pseudo-charge contained in a closed domain is $mQ$ where $m$ is the difference between the number of walls crossing the path inwards and outwards (Supplementary Note). All pseudo defects have polarity, defined by the direction of the outgoing wall for startpoints (Fig. 3) and the direction of the incoming wall for branchpoints and endpoints (Figs. 3 and 3).
The topological pseudo-charge is preserved by the dynamics of the director. Pseudo-defects with opposite charges are created in pairs and annihilated in pairs (Figs. 3, 3 and 3), similarly to actual nematic defects. A wall endpoint cannot merge with a preexisting wall. Instead, it can connect with a startpoint ahead of it (the inverse process of Fig. 3) or, through wall shrinkage, it can annihilate with the nearest negative pseudo-defect it is connected to (the inverse processes of Figs. 3 and 3). Notably, this nearest connected point does not need to be its birth pair and may be quite a long distance away, as we often observe in the arrested regime. In addition to pair birth and annihilation, negative pseudo-defects can transition from one to the other while conserving the pseudo-charge (Fig. 3 and its inverse process). In another pseudo-topological transition, connected branchpoints can exchange branches (Fig. 3). This process is akin to a T1 transition in foams and tissues, except that our lines are directed, and branching nodes are restricted to having one incoming wall and two outgoing walls. As a result, T2 transitions are not possible.
@@
Figure 4: Motifs of arrested wall networks and experimental snapshots of microtubule-kinesin active nematics. In all diagrams, lines, nodes and colors are as defined in Fig. 3. 44, Basic network motifs. The anchoring motif (4) is made of an endpoint and branchpoint that meet head-on, with the endpoint trapped between the two outgoing walls of the branchpoint. In motif (4) the endpoint meets the branchpoint from one of its sides, i.e. between the incoming wall and an outgoing wall. The dashed line traces a weak distortion, indicating that the wall associated with the endpoint tends to align its direction with the outgoing wall on the opposite side of the branchpoint. The motifs (4,4) involve a single pseudo-defect interacting with a bare wall. These, along with (4), do not follow the tendency to have strictly antiparallel walls. 4, Composite motif schematic (left) and one formed spontaneously in a simulation (right). The stream plot on the right represents the flow, with black indicating maximal $|\bm{v}|$ and full transparency indicating $|\bm{v}|=0$. The gray background is the line-integral-convolution representation of the nematic director $\bm{n}$. Parameter values and color legend for splay and bend distortions are as in Fig. 1. 44, Raw fluorescence images from experiments (left panels) and overlaid schematic drawings (right panels) depicting domain walls, pseudo-defects and actual ±$\pm 1/2$ defects in white. 4, Taken from a movie in 49 (courtesy of Guillaume Duclos), which shows the evolution of the microtubule-based nematic following the bending instability of the aligned state. 4, Taken from a movie in 50 (courtesy of Pau Guillamat), which shows a turbulent transient with all types of pseudo-defects and actual nematic defects. Note how walls may also originate from true $+1/2$ defects and be absorbed by true $-1/2$ defects.
Motifs of arrested wall networks
@@
In the arrested turbulence regime (with $S\nu<0$), the processes of wall folding, branching, extension, and arrest produce a directed tree-like network of self-avoiding walls (Fig. 2). This network does not form closed loops in a simply-connected domain because such a loop would enclose a topological charge of 1, implying actual nematic defects, which our system forbids.
@@
We find that endpoints propagate rapidly in the direction of their polarity, until they are blocked by other walls. Startpoints, on the other hand, move quickly in the opposite direction of their polarity until their wall dissolves or they merge with another wall, transitioning into a branchpoint. Unlike endpoints and startpoints, branchpoints do not move in the direction parallel or antiparallel to their polarity. Rather, when they are created (Fig. 3), they tend to deflect only slightly in the direction of the newly formed branch (Fig. 2). Hence, in the sense of their mobility, endpoints are akin to actual $+1/2$ nematic defects and branchpoints are similar to $-1/2$ defects, which are not mobile.
@@
Ultimately, the pseudo-defects—primarily endpoints and branchpoints—maintain a steady distance from neighbouring walls and organize into motifs like those shown in Figs. 4 to 4. A frequent motif, which we call the anchoring motif (Fig. 4), is the one discussed previously in relation to the stagnation points in Fig. 1. We postulate that this structure might be particularly stable because neighbouring walls are anti-parallel, as promoted by the spontaneous-flow instability. In contrast, the other basic motifs (Figs. 4 to 4) feature some neighbouring walls that are parallel. For smaller system sizes, closer to the transition to chaos, we observe that only the anchoring motifs survive in the long time limit. For large system sizes, a variety of composite structures combining all types of motifs can emerge and persist for long times. One such structure is shown in Fig. 4.
@@
We identified bending domain walls, pseudo-defects and our motifs in images and videos of microtubule-based active nematics 51; 50; 15; 49; 52 (Figs. 4 to 4). In Fig. 4, the spontaneous-flow instability leads to the growth of bending domain walls and the emergence of pseudo-defects. These tend to organize in interlaced anchoring motifs, which are also frequently observed in our simulations. In Fig. 4, pseudo-defects are seen to coexist with actual nematic defects (white). Since the latter may also act as origin points or termination points of walls, there may be an interesting interplay between both types of structures in organising the dynamics of the system. It is important to note that in these experimental examples the system does not maintain an arrested state. We expand on this aspect further in the final discussion.
@@
Figure 5: Ageing of the arrested wall network. 5,5, Skeleton of the domain walls (black) with startpoints, branchpoints and endpoints (green, blue and red triangular nodes) at an early time (5) and a late time (5). The detection of the network skeleton and its nodes is described in Methods .3, Supplementary Fig. 4. 5, Evolution of the number of startpoints (green), branchpoints (blue) and endpoints (red). In the initial transient, sequential ‘zig-zag’ instabilities result in the proliferation of both branchpoints and endpoints. Once the wall pattern establishes a wavelength, the system ages slowly as some endpoints retract and annihilate with their connected branchpoint, while others extend (Supplementary Movie 5). Throughout the simulation, there are frequent transitions between branchpoints and startpoints, though the number of startpoints remains low. Additionally, the detection algorithm is not perfect, occasionally misidentifying endpoints or branchpoints as startpoints and vice versa. Parameter values were set to $R=1$, $S=1$, $\nu=-0.9$, and $A=3.2\times 10^{5}$
Ageing dynamics
The dynamics of active nematics are not relaxational, and the arrested patterns reported here co-exist with a background of weak large-scale chaotic flows. Nevertheless, these dynamics share interesting similarities with relaxational systems that exhibit frustration. In particular, trapping motifs (Figs. 4 to 4) might promote arrest similarly to how cages and locally-favoured structures promote arrest in gels and glasses 53; 54; 55. Additionally, the wall network undergoes a slow relaxation that is reminiscent of the ageing of gels and glasses. The background turbulent-like flows produce small low-$q$ fluctuations that slowly evolve the pattern. This includes jittering the walls and inducing some pseudo-topological transitions in the network; primarily those depicted in Figs. 3, 3 and 3 and their inverse processes. Crucially, these represent a highly restrictive subset of transitions compared to the full range of processes enabled when startpoints are more prominently present. Through this fluctuating process, we find that the number of pseudo-defects gradually decreases (Fig. 5, Supplementary Movie 5), mostly via annihilation of connected branchpoint-endpoint pairs. As some walls shrink and disappear from the network, others extend, thereby maintaining the selected wavelength. At long times ($\sim 10^{3}\,\tau_{\text{a}}$), the evolution slows down and the number of pseudo-defects seems to approach a steady state (Fig. 5).
@@
Figure 6: Unicursal labyrinths. 6, Two examples of arrested tree-like patterns formed by domain walls, with $S\nu=-0.9$ (left) and $S\nu=-1.1$ (right). 6, Processed negatives of (6) showing how domain walls (black) enclose a small and even number of unicursal labyrinths (4 connected regions on the left, 2 on the right) that span the entire system (Methods .3, Supplementary Fig. 5). The walls separating different labyrinths are self-connected across the periodic boundaries, unlike all other branches which feature endpoints. 6, The mythical labyrinth of Crete, shown on a silver coin from around 400 BC (by AlMare, reproduced from Wikimedia Commons, creative commons license CC BY-SA 3.0). 6, The labyrinth in the Chartres cathedral (courtesy of Père Emmanuel Blondeau and the secretariat de la Cathédrale Notre-Dame de Chartres). 6, Meander pattern known as the Greek key encircling a box (pyxis) from around 850 BC (reproduced from the Metropolitan Museum of Art). 6, Out-of-plain magnetization domains of a ferrimagnetic garnet film, with black and white regions indicating opposite magnetizations. The amplitude of an external magnetic field favouring the black domain is gradually reduced from left to right, resulting in growth by folding of the white unicursal domain. Frames are desaturated snapshots from a video courtesy of Pietro Tierno.
The tree-like wall network encloses unicursal labyrinths
@@
When there are no startpoints, the gaps between the walls form unicursal labyrinths of a type known as meanders, defined by having a single path with neither bifurcations nor dead ends. A typical wall network encloses a small and even number of such labyrinths, with each one closing on itself through the periodic boundaries, i.e., through the hole of the torus, which is not a simply-connected domain (Figs. 6 and 6). On the other hand, the wall network bounding these unicursal paths forms a maze, defined as having multiple bifurcations and dead ends, corresponding to our branchpoints and endpoints. The specific walls separating the different unicursal labyrinths are initially generated by the spontaneous-flow instability. There are as many of these primary walls as there are unicursal labyrinths, and they also close on themselves through the periodic boundaries. Since primary walls are the only paths on the network that avoid dead ends, they can be thought of as the solutions to the maze.
@@
Paintings, rock-carvings, ornaments and monuments depicting unicursal labyrinths predate recorded history and have remained common in artistic and spiritual designs 56. Renowned examples include the Cretan labyrinth of Greek mythology (Fig. 6), and the labyrinth decorating the floor of Chartres cathedral (Fig. 6). However, unlike ours, these and most other Labyrinth designs feature a dead end in the center. On the other hand, meanders (Fig. 6) may close on themselves through the hole of the torus. In pattern-forming systems, unicursal labyrinths are rare. They were obtained under special conditions in simulations of a modified Swift-Hohenberg equation 57, and in experiments with ferrimagnetic garnet films in specific experimental protocols 58; 59; 60 (Fig. 6). In our case, they result from an emerging topology of domain wall networks and they occur spontaneously and generically in the arrested regime.
Discussion
In this work, we showed that flow alignment—the tendency of liquid crystals to reorient under shear—dramatically influences the spatiotemporal structure of defect-free active nematic turbulence (Fig. 1). For contractile flow-aligning active nematics, the chaos is characterized by fragmented nematic patterns, reduced total Frank free energy, and strong self-similar flows across a broad range of scales. We associate this strong large-scale turbulence to the underlying metastability of uniform and wall-laden states 37. On the other hand, despite their chaotic dynamics, we find that extensile aligning active nematics self-organize into an arrested tree-like pattern of nematic domain walls, with the flow strongly channeled along them (Figs. 1 and 1). This finding reveals a mechanism of pattern formation mediated by active turbulence, complementing those found in other studies 61; 62; 63. In particular, our patterns enclose unicursal labyrinths (Fig. 6), which do not appear spontaneously in other pattern-forming systems.
The labyrinths in Fig. 6 span the entire system, but how far can they possibly extend? Further increasing the system size widens the range of chaotic self-similar flows, identified by the $q^{-1}$ scaling of the velocity power spectrum (Fig. 1, blue). These flows become stronger at larger scales, and hence they could potentially disrupt the arrested wall pattern for sufficiently large systems. At the scale where chaotic large-scale flows become as strong as the wall streams, which sustain the domain walls, the tree-like network could lose its connectivity and fluctuate more strongly. In fact, for the system size studied here, this competition between wall streams and chaotic large-scale flows is observed for $\nu=0$ (Fig. 1, black). In this case, we observe a wall network, locally similar to the arrested ($S\nu=-1.1$) pattern, but sporadically fractured in random positions (Supplementary Fig. 1, Supplementary Movie 3). In line with our hypothesis, the fractured areas where the wall pattern dissolves (dark regions in the Frank free energy density) appear to coincide with hot spots of large-scale flows. The wall pattern is subsequently restored at fractured regions because large jets are not persistent, as shown by the low correlation times of low-$q$ flow in Fig. 1, dashed black.
The arrested state does not show long-range correlations in the nematic order parameter (Supplementary Fig. 2, blue). This is reminiscent of some passive systems, such as spin ice, that feature frustration in the absence of long-range geometrical correlations64. Similarly, our pattern features short-range disorder and is yet connected over long distances. Therefore, we propose that the unicursal labyrinths exhibit a form of long-range topological order stemming from emergent connectivity rules associated with the folding, branching, extension and self-avoidance of the domain walls.
In the fully-developed turbulent state, both the arrested and non-arrested regimes manifest a selected wavelength. This wavelength is significantly greater than the active length $l_{\text{a}}$ (Fig. 1), and it also exceeds the critical wavelength of the spontaneous-flow instability. The wavelength selection mechanism is inherently nonlinear and two-dimensional; it is based on a balance between the coarsening of local striped patterns, which tends to enlarge the characteristic wavelength (Fig. 2), and the bending and folding of walls that tends to reduce it (Fig. 2). A deeper understaning of the selection mechanism and its dependence on parameters remains an open question.
Whereas our model is defect-free, most experiments and simulations in active nematics have been conducted in defect-laden regimes. Yet, the physics of the two regimes may intersect, perhaps similarly to the crossovers found in polar fluids 65; 66. Actual nematic defects tend to nucleate along nematic domain walls, thereby dissolving them 43; 44; 45; 34. Additionally, nematic constituents such as microtubules are depleted at defects (Fig. 4), which lowers their energy cost and favours their nucleation67. The frequent dissolution of domain walls by defects explains why, in the instances where our motifs can be seen (Figs. 4 to 4), they do not last for long. We therefore propose that, in addition to their well-known role as drivers of turbulence by generating vorticity 32, defects also play an important role as enablers of stronger turbulent flows by breaking the potential grid-lock, i.e., preventing dynamical arrest.
Directly testing our predictions for defect-free active nematics presents an experimental challenge. This would require searching for active materials for which the nucleation of defect pairs were significantly more costly than in the experimental realizations studied thus far. Defect-free active turbulence should be achievable for sufficiently large system sizes $L$, ensuring large dimensionless activity $A=(L/\ell_{\text{a}})^{2}\gg 1$, while keeping the defect-core size $a$ small enough ($\ell_{\text{a}}\gg a$) to preclude defect nucleation.
In conclusion, our study of defect-free active nematic turbulence highlights the crucial role of topology beyond the existence of topological defects. We argue that the connectivity of domain walls, the existence of pseudo-defects, and their interactions are key to understanding the dynamics of active nematics, from strong large-scale turbulence to dynamical arrest. Furthermore, when true defects are present, their interplay with pseudo-defects likely adds another intriguing layer of complexity. Our results open new avenues for experimental analysis and design, particularly in tracing the intricate web of domain walls, their orientation, and their nodes.
Acknowledgments
Funding from the Spanish Ministry for Science and Innovation MICCINN/FEDER (PID2022-137713NB-C22 to IL and JC) and the Generalitat de Catalunya (AGAUR SGR-2021-00450 to IL and JC and ‘ICREA Academia’ award to JC) is acknowledged. IL acknowledges support from the Flatiron Instutute, part of the Simons Foundation. RA thanks the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the Programme New Statistical Physics in Living Matter, supported by EPSRC Grant Number EP/R014604/1, when work on this paper was undertaken.
Author contributions
IL performed all the simulations and analyses. JC conceived and supervised the research. IL and JC developed the theory. All authors discussed the results and wrote the manuscript.
Competing interests
The authors declare no competing interests.
Data and code availability
The data and code supporting the findings of this study are available from the corresponding author, IL, upon reasonable request.
Supplementary Figures
@@ Acknowledgments
Supplementary Figure 1: Active nematic turbulence in the absence of flow-alignment. Snapshots from a simulation of defect-free active nematic turbulence with parameter values $R=1$, $\nu=0$, and $A=3.2\times 10^{5}$. We chose a contractile stress ($S=-1$), but the results are exactly equivalent in the extensile case ($S=+1$) when the director is rotated ($\theta\to\theta\pm\pi/2$). 1, The flow field; black curves are streamlines, and the color indicates the speed (Supplementary Movie 3, left). 1, Frank free energy density $\sim|\nabla\theta|^{2}$, with high-intensity lines corresponding to nematic domain walls (Supplementary Movie 3, right). Note how sparse peaks in the flow (red regions in 1) tend to coincide with perturbed wall patterns (dark regions in 1). 11, Zooms in the contractile (1) and extensile (1) cases highlighting the type of nematic distortion as well as the interplay between nematic walls and flows. The gray-scale background is the line integral convolution representation of the director field $\bm{n}$. Magenta and cyan intensities respectively represent splay $(\nabla\cdot\bm{n})^{2}$ and bend |×$|\nabla\times\bm{n}|^{2}$ contributions to the Frank energy density. The black arrows represent the flow field $\bm{v}$.
Supplementary Figure 2: Power spectrum of the order parameter in active nematic turbulence. Shown is the power spectrum of the nematic tensor, $\hat{q}_{\alpha\beta}=n_{\alpha}n_{\beta}-1/2\,\delta_{\alpha\beta}$, in the contractile flow-aligning ($S\nu=+1.1$, red), extensile flow-aligning ($S\nu=-1.1$, blue), and non-aligning ($\nu=0$, black) cases. For the definition of $W(q)$ see Methods .2. Lines represent a smoothed (Gaussian) interpolation of the computed data points. Parameter values are as in Fig. 1 and Supplementary Fig. 1. For arrested turbulence (blue), this spectrum features a narrow peak with its associated harmonics. For strong large-scale turbulence (red), such peaks are washed out. In all regimes, low-$q$ correlations are essentially absent.

### v3

@@ IV.1 Wall nodes define pseudo-defects
These recurring node structures arise naturally in our simulations and serve as an effective framework for characterizing defect-free dynamics. We note that a node with two incoming walls and one outgoing wall (a mergepoint) is only possible by combining different wall types (Fig. S235). Additionally, we observethat startpoints are very common in the non-arrested (contractile aligning) regime but become increasingly rare over time in the arrested (extensile aligning) regime.
[…]
The topological pseudo-charge is preserved by the dynamics of the director. Pseudo-defects with opposite charges are created in pairs and annihilated in pairs (Figs. 3, 3 and 3, Animations S1 to S3), similarly to actual nematic defects. A wall endpoint cannot merge with a preexisting wall. Instead, it can connect with a startpoint ahead of it (the inverse process of Fig. 3, Animation S2) or, through wall shrinkage, it can annihilate with the nearest negative pseudo-defect it is connected to (the inverse processes of Figs. 3 and 3, Animations S1 and S3). Notably, this nearest connected point does not need to be its birth pair and may be quite a long distance away, as we often observe in the arrested regime. In addition to pair birth and annihilation, negative pseudo-defects can transition from one to the other while conserving the pseudo-charge (Fig. 3 and its inverse process, Animation S4). In another pseudo-topological transition, connected branchpoints can exchange branches (Fig. 3, Animation S5). This process is akin to a T1 transition in foams and tissues, except that our lines are directed, and branching nodes are restricted to having one incoming wall and two outgoing walls. As a result, T2 transitions are not possible.
@@
Figure 4: Motifs of arrested wall networks and experimental snapshots of microtubule-kinesin active nematics. In all diagrams, lines, nodes and colors are as defined in Fig. 3. 44, Basic network motifs. The anchoring motif (4) is made of an endpoint and branchpoint that meet head-on, with the endpoint trapped between the two outgoing walls of the branchpoint. In motif (4) the endpoint meets the branchpoint from one of its sides, i.e. between the incoming wall and an outgoing wall. The dashed line traces a weak distortion, indicating that the wall associated with the endpoint tends to align its direction with the outgoing wall on the opposite side of the branchpoint. The motifs (4,4) involve a single pseudo-defect interacting with a bare wall. These, along with (4), do not follow the tendency to have strictly antiparallel walls. 4, Composite motif schematic (left) and one formed spontaneously in a simulation (right). The stream plot on the right represents the flow, with black indicating maximal $|\bm{v}|$ and full transparency indicating $|\bm{v}|=0$. The gray background is the line-integral-convolution representation of the nematic director $\bm{n}$. Parameter values and color legend for splay and bend distortions are as in Fig. 1. 44, Raw fluorescence images from experiments (left panels) and overlaid schematic drawings (right panels) depicting domain walls, pseudo-defects and actual $\pm 1/2$ defects in white. 4, Taken from a movie in 50 (courtesy of Guillaume Duclos), which shows the evolution of the microtubule-based nematic following the bending instability of the aligned state. 4, Taken from a movie in 51 (courtesy of Pau Guillamat), which shows a turbulent transient with all types of pseudo-defects and actual nematic defects. Note how walls may also originate from true +1/2+1/2 defects and be absorbed by true 1/2-1/2 defects.
IV.2 Network motifs
@@
In the arrested turbulence regime (extensile rods with $S\nu<0$), the processes of wall folding, branching, extension, and arrest produce a directed tree-like network of self-avoiding walls (Fig. 2). This network does not form closed loops in a simply-connected domain because such a loop would enclose a topological charge of 1, implying actual nematic defects, which our system forbids.
@@
We find that endpoints propagate rapidly in the direction of their polarity, until they are blocked by other walls. Startpoints, on the other hand, move quickly in the opposite direction of their polarity until their wall dissolves or they merge with another wall, transitioning into a branchpoint. Unlike endpoints and startpoints, branchpoints do not move in the direction parallel or antiparallel to their polarity. Rather, when they are created (Fig. 3), they tend to deflect only slightly in the direction of the newly formed branch (Fig. 2). Hence, in the sense of their mobility, endpoints are akin to actual +1/2+1/2 nematic defects and branchpoints are similar to 1/2-1/2 defects, which are not mobile. Despite these similarities, we emphasize that actual defects generate fundamentally different dynamics, likely due to fewer constraints on their mobility and the transitions they can undergo. This contrast becomes evident in the final section, where we perform full Q-tensor simulations that permit the nucleation of true topological defects.
@@
Ultimately, the pseudo-defects—primarily endpoints and branchpoints (in arrested states)—maintain a steady distance from neighbouring walls and organize into motifs like those shown in Figs. 4 to 4. A frequent motif, which we call the anchoring motif (Fig. 4), is the one discussed previously in relation to the stagnation points in Fig. 1. We postulate that this structure might be particularly stable because neighbouring walls are anti-parallel, as promoted by the spontaneous-flow instability. In contrast, the other basic motifs (Figs. 4 to 4) feature some neighbouring walls that are parallel. For large system sizes, a variety of composite structures combining all types of motifs can emerge and persist for long times (see Movie S2). One such structure is shown in Fig. 4.
@@
We identified bending domain walls, pseudo-defects and our motifs in images and videos of microtubule-based active nematics 52; 51; 15; 50; 53; 54 (Figs. 4 to 4). In Fig. 4, the spontaneous-flow instability leads to the growth of bending domain walls and the emergence of pseudo-defects. These tend to organize in interlaced anchoring motifs, which are also frequently observed in our simulations. In Fig. 4, pseudo-defects are seen to coexist with actual nematic defects (white). Since the latter may also act as origin points or termination points of walls, there may be an interesting interplay between both types of structures in organizing the dynamics of the system. We take initial steps toward exploring this interaction in the final section, where we analyze the full Q-tensor model. We further note that a very recent study 54 observed that a series of folding events generates an extensive network of bending domain walls, where both types of nodes (pseudo-defects and true defects) are clearly present. Importantly, in these experimental examples, the system does not maintain an arrested state. We expand on this aspect further in the final discussion.
@@
Figure 5: Ageing of the arrested wall network. $5,$ Skeleton of the domain walls (black) with startpoints, branchpoints and endpoints (green, blue and red triangular nodes) at an early time (5) and a late time (5). The detection of the network skeleton and its nodes is described in Appendix E, Fig. 13. 5, Evolution of the number of startpoints (green), branchpoints (blue) and endpoints (red). In the initial transient, sequential ‘zig-zag’ instabilities result in the proliferation of both branchpoints and endpoints. Once the wall pattern establishes a wavelength, the system ages slowly as some endpoints retract and annihilate with their connected branchpoint, while others extend (Movie S5). Throughout the simulation, there are frequent transitions between branchpoints and startpoints, though the number of startpoints remains low. Additionally, the detection algorithm is not perfect, occasionally misidentifying endpoints or branchpoints as startpoints and vice versa. Parameter values were set to $R=1$, $S=1$, $\nu=-0.9$, and $A=3.2\times 10^{5}$.
IV.3 Ageing-like dynamics
The dynamics of active nematics are not relaxational, and the arrested patterns emerging in the extensile case co-exist with a background of weak large-scale chaotic flows. Nevertheless, these dynamics share interesting similarities with relaxational systems that exhibit frustration. In particular, trapping motifs (Figs. 4 to 4) might promote arrest similarly to how cages and locally-favoured structures promote arrest in gels and glasses 55; 56; 57. After an initial transient in which sequential branching establishes the network, we find that the wall network undergoes a slow relaxation that is reminiscent of the ageing of gels and glasses. The background turbulent-like flows produce small low-$q$ fluctuations that slowly evolve the pattern. This includes jittering the walls and inducing some pseudo-topological transitions in the network; primarily those depicted in Figs. 3, 3 and 3 and their inverse processes. Crucially, these represent a highly restrictive subset of transitions compared to the full range of processes enabled when startpoints are more prominently present. Through this fluctuating process, we find that the number of pseudo-defects gradually decreases (Fig. 5, Movie S5), mostly via annihilation of connected branchpoint-endpoint pairs. As some walls shrink and disappear from the network, others extend, thereby maintaining the selected wavelength. At long times ($\sim 10^{3}\,\tau_{\text{a}}$), the evolution slows down and the number of pseudo-defects seems to approach a steady state (Fig. 5). This relaxation occurs on timescales of hundreds of $\tau_{\text{a}}$ and appears to be a specific feature of arrested wall networks. It is absent in the contractile aligning regime, and in the presence of true topological defects—as shown in the final section.
@@ IV.3 Ageing-like dynamics
Figure 6: Unicursal labyrinths. 6, Two examples of arrested tree-like patterns formed by domain walls, with $S\nu=-0.9$ (left) and $S\nu=-1.1$ (right). 6, Processed negatives of (6) showing how domain walls (black) enclose a small and even number of unicursal labyrinths (4 connected regions on the left, 2 on the right) that span the entire system (Appendix E, Fig. 14). The walls separating different labyrinths are self-connected across the periodic boundaries, unlike all other branches which feature endpoints. 6, The mythical labyrinth of Crete, shown on a silver coin from around 400 BC (by AlMare, reproduced from Wikimedia Commons, creative commons license $C$ BY-SA 3.0). 6, The labyrinth in the Chartres cathedral (courtesy of Père Emmanuel Blondeau and the secretariat de la Cathédrale Notre-Dame de Chartres). 6, Meander pattern known as the Greek key encircling a box (pyxis) from around 850 BC (reproduced from the Metropolitan Museum of Art). 6, Out-of-plain magnetization domains of a ferrimagnetic garnet film, with black and white regions indicating opposite magnetizations. The amplitude of an external magnetic field favouring the black domain is gradually reduced from left to right, resulting in growth by folding of the white unicursal domain. Frames are desaturated snapshots from a video courtesy of Pietro Tierno.
IV.4 Unicursal labyrinths
@@
When there are no startpoints, which is typical in the extensile aligning regime, the gaps between the walls form unicursal labyrinths of a type known as meanders, defined by having a single path with neither bifurcations nor dead ends. A typical wall network encloses a small and even number of such labyrinths, with each one closing on itself through the periodic boundaries, i.e., through the hole of the torus, which is not a simply-connected domain (Figs. 6 and 6). On the other hand, the wall network bounding these unicursal paths forms a maze, defined as having multiple bifurcations and dead ends, corresponding to our branchpoints and endpoints. The specific walls separating the different unicursal labyrinths are initially generated by the spontaneous-flow instability. There are as many of these primary walls as there are unicursal labyrinths, and they also close on themselves through the periodic boundaries. Since primary walls are the only paths on the network that avoid dead ends, they can be thought of as the solutions to the maze.
@@
Paintings, rock-carvings, ornaments and monuments depicting unicursal labyrinths predate recorded history and have remained common in artistic and spiritual designs 58. Renowned examples include the Cretan labyrinth of Greek mythology (Fig. 6), and the labyrinth decorating the floor of Chartres cathedral (Fig. 6). However, unlike ours, these and most other Labyrinth designs feature a dead end in the center. On the other hand, meanders (Fig. 6) may close on themselves through the hole of the torus. In pattern-forming systems, unicursal labyrinths are rare. They were obtained under special conditions in simulations of a modified Swift-Hohenberg equation 59, and in experiments with ferrimagnetic garnet films in specific experimental protocols 60; 61; 62 (Fig. 6). In our case, they result from an emerging topology of domain wall networks and they occur spontaneously and generically in the arrested regime.
@@
V Unconstrained Q-tensor model
To place our results in a broader context, we now turn to the unconstrained active Q-tensor model, based on the Landau-de Gennes framework, which allows for the nucleation of true $\pm 1/2$ defect pairs. This enables us to (i) delineate the regime in which our defect-free findings remain applicable, and (ii) explore the transition to the more familiar defect-laden behavior that arises when those conditions are no longer met. As in many previous studies, e.g. 63; 64; 65; 66; 67; 34; 68, we use the governing equations based on the Beris-Edwards model 69 for dimension $d=2$ and $\text{Re}=0$ with the addition of the active nematic stress. These equations, expressed in our notation, together with details of the numerical scheme used to evolve $\mathbf{Q}$, are provided in Appendix D. For our discussion here, we define the nematic tensor in terms of the scalar order parameter $s$, quantifying local nematic alignment, and a unit director $\bm{n}$, as
$Q_{\alpha\beta}=s\left(n_{\alpha}n_{\beta}-\frac{1}{2}\delta_{\alpha\beta}\right).$
With respect to our constrained director-based formulation presented earlier, the key difference lies in the system’s free energy: